OpenStudy (anonymous):

Use logarithmic differentiation to find the derivative of the function. y= x^(4cosx)

OpenStudy (zehanz):

Take ln on both sides. Use lna^b = b*ln a Differentiate left and right Multiply with y

OpenStudy (anonymous):

This is what I got from differentiating on both sides: (1/y)*y'=-4sinx*lnx+(1/x)*4cosx

OpenStudy (anonymous):

good, now just go move your y to the numerator on the other side and change it back to what you started out with, which was y=x^(4cosx)

OpenStudy (zehanz):

OK, so now multiply left and right with y to get y'

OpenStudy (anonymous):

Yeah, what ZeHanz said.

OpenStudy (zehanz):

And be sure to replace y with x^(4cosx), like @calmat01 said :)

OpenStudy (anonymous):

lol..Thanks, @ZeHanz

OpenStudy (anonymous):

I got y'=-4sin(x)*lnx+(1/x)*4cos(x)*x^(4cosx)

OpenStudy (anonymous):

Make sure you put the brackets around the right terms. To be on the safe side, I would put the x^(4cosx) first and then put everything after it in one big bracket.

OpenStudy (anonymous):

so it would be x^(4cosx)*[-4sin(x)*lnx+(1/x)*4cosx]?

OpenStudy (anonymous):

There ya go.

OpenStudy (zehanz):

Yes, though you could write 1/x*4cosx as \[\frac{ 4\cos x }{ x }\]

OpenStudy (anonymous):

you could probably simplify that 1/x)..lol

OpenStudy (anonymous):

lol Ya beat me to it, @ZeHanz

OpenStudy (zehanz):

Yes, it's fun isn't it!

OpenStudy (anonymous):

could the answer also be 4x^(4cos(x))*[cosx/(x)-ln(x)*sin(x)]?

OpenStudy (anonymous):

Yes, it could. By factoring out a 4 from both terms.

OpenStudy (anonymous):

awesome! thank you both for helping me through the problem :)

OpenStudy (anonymous):

Np. Glad to be of help.

OpenStudy (zehanz):

YW!